# Identifying companies and financial actors exposed to marine tipping points

### v1

@@ Introduction
Increasing sea surface temperature can destabilize upwellings – currents that bring nutrient-rich water to the ocean’s surface – and affect some of the world’s most productive fisheries11. While insufficient nutrients can reduce primary productivity and cascade up through marine food webs, an excess can cause coastal eutrophication and hypoxia6, with detrimental consequences to fish stocks. Climate change is expected to reduce fish biomass by 3 to 22% in tropical and temperate areas by the end of the century12. Overfishing and nutrient pollution from agriculture and urban waste can exacerbate this impact. A combination of oceanographic and climatic factors heightens the risk of regime shifts, putting both ecosystems and human livelihoods at risk2,13. However, proactive management strategies that account for these risks could help mitigate exposure to tipping points2.
@@
Companies and financial institutions are key actors that contribute to the extraction of marine resources, while also being affected economically by changes in marine ecosystems. About 35.4% of fish stocks globally are currently depleted or overexploited14, with only 13 companies controlling around 16% of wild seafood capture, and just five companies responsible for 48% of farmed salmon production15,16. Fishing companies contribute to ecological degradation through harmful practices like overfishing17, destructive fishing methods18, and non-compliance with regulations19,20. By removing biomass from the ocean, fishing has historically prevented the sequestration of 21.8M tons of carbon14,21. In addition, the industry’s return on investment is becoming increasingly risky. Modern fishing fleets travel twice as far but catch only one-third of what they used to per kilometre travelled22, with 54% of high seas fishing would be unprofitable without public subsidies14,23.
@@
Figure 1: Areas at risk of tipping points. We computed extreme and severe heatwave events based on historical records of sea surface temperature (A, B). Areas in colour are places where the probability is non-zero, expressed as the mean number of days spent on a heatwave event. Ocean colour data was used to compute early warning signals of critical transitions (C). Note that there are no warnings computed for higher latitudes due to missing data. For the analysis of areas prone to marine regime shifts, we kept areas where the probability of extreme and severe heatwaves is non-zero, or where at least three early warning signals were found (see Methods).
@@
However, it should be in the best interest of companies and their shareholders to act as stewards of the seas and maintain a productive and healthy ocean24,25. Financial actors – such as banks, asset managers, and pension funds – could play an important role in conservation by deciding not to finance economic activities that alter the ecological integrity of the biosphere26–29. For instance, a small set of financial actors have an increasing influence over climate stability through their equity holdings in companies operating in the Amazon and boreal forests26. Likewise, large asset managers and other financial institutions invest in companies operating in areas vulnerable to the emergence of infectious diseases27.
@@
This leaves them exposed to the impacts of their investments30. Climate-change-driven extreme weather events are expected to cost up to USD 14 trillion annually by 210031, directly threatening the solvency of insurers and reinsurers32. Investors and asset managers face climate exposure being at risk of stranded assets, changes in commodity prices, scarcity of resources, disruption of supply chains, and infrastructure damages33. In response, financial actors are increasingly expected to disclose their risks through initiatives such as the Task Forces on Climate- and Nature-Related Financial Disclosures (TCFD and TNFD, respectively) or the EU non-financial reporting directive.
@@
Analyses of financial exposure to biodiversity loss and ecosystem degradation are mostly focused on terrestrial ecosystems, with little research addressing ocean-related financial risks34. Yet, fishing companies, their shareholders, and the countries they operate in are exposed to potentially irreversible changes and consequent declines in ecosystem services, with consequences on the economy, investments and livelihoods. However, being exposed to regime shifts also means being in a unique position to take action, avoid tipping points, and contribute to sustainable transformations. In this paper, we asked who are these corporate and financial actors exposed to marine tipping points?
@@ Methods
Areas at risk: We identified areas of the world vulnerable to marine regime shifts. We approximated these areas as places currently under the strong influence of marine heatwaves or already showing symptoms of resilience loss. Frölicher et al.8 defined heatwaves as events when sea surface temperature exceeds the 99th percentile with respect to daily satellite observations for a 30-year period. Hobday et al.35 further classified heatwave categories based on the 90th percentile, where moderate heatwaves are between 1x and 2x the 90th percentile, strong heatwaves between 2x and 3x, severe heatwaves between 3x and 4x, and extreme heatwaves above 4 times the 90th percentile. We computed the probability of severe and extreme heat waves based on daily sea surface temperature observations from the US National Oceanic and Atmospheric Administration (NOAA) using the optimum interpolation data from 1981 to 2022 at 0.25 resolution (Fig 1A). While several impacts of heatwaves have been documented in ecosystems and the economy10,35, a recent global assessment showed no systematic effect on fish biomass36. However, heatwaves by themselves are seldom the sole cause of fish populations decline, it is often the compound effect of temperature with other pressures such as low primary productivity or low oxygen zones13.
@@
Thus we complemented heat waves detection with proxies stability loss in primary productivity, which are in turn driven by oxygen depletion, salinity, and the mixing of the upper layer of the ocean6,7,37. We used a geographically explicit mapping of marine ecosystems showing symptoms of resilience loss38. Resilience is the ability of a system to withstand disturbances without losing its functions and identity39. When systems lose resilience and are close to tipping points, they leave statistical signatures in time series known as early warning signals40. The map was produced by computing temporal autocorrelation, standard deviation, skewness, kurtosis and fractal dimension on Chlorophyll-A concentration, a proxy of marine primary productivity, using data from the European Space Agency (weekly observations at 0.25 resolution from 1998-2018)41. The data are limited by the coverage of the ocean colour data, excluding the Arctic and Antarctic oceans38. For this paper we used areas with at least 3 early warnings, since two of them are suitable for bifurcation tipping (autocorrelation and standard deviation), two of them are suitable for flickering and stochastic transitions (skewness and kurtosis), and the fractal dimension is robust against heteroskedasticity. At least three early warnings give us the chance to pick up more robust signals of the risk of tipping taking into account different ways of reaching the transition (Fig 1).
@@
Fishing pressure: We identified fishing vessels operating in pixels showing a high probability of heatwaves or symptoms of resilience loss. We used the dataset published by Kroodsma et al.42 who identified >70,000 industrial fishing vessels in the size range from 6 to 146m in length, thanks to an automatic identification system (AIS). Vessels need to broadcast their position on the AIS to avoid potential collisions, so the data offers individual vessel trajectories from 2012-2020 (fishing effort dataset from Global Fish Watch version 2.0). We extracted vessels whose fishing activities were recorded in areas with high probabilities of heatwaves (severe or extreme events) or high probability of regime shifts. We only include pixels with at least three different early warning signals of resilience loss, that way we include multiple routes of tipping (e.g. noise-induced, flickering, not only the traditional slowing down), and keep pixels for which multiple lines of evidence suggest loss of resilience.
@@
Owners and shareholders: We identified the individual and company beneficial owners of the vessels using the maritime mobile service identity to match the Global Fishing Watch data with vessel owners in the Lloyds database (www.seasearcher.com). We matched 13,537 records, but not all of them had available ownership information. We also used the dataset curated by Carmine et al.43 where vessels fishing on the high seas were manually annotated. We added companies whose vessels matched our selection of vessels fishing in areas with marine tipping point risk. To identify shareholders and global ultimate owners of the companies responsible for fishing activities, we used Orbis, a database of companies around the world with coverage of > 489 million companies. Where possible, we extracted data about their shareholder structure (e.g. direct and total ownership, shareholder type), annual revenues, number of employees, and shareholder’s countries of operation. Unfortunately, Orbis and Lloyds are not open access data sources and their terms of service forbids sharing the raw accessed data with third parties. Both of them were accessed through research subscriptions by the home institution.
@@
Network models: We constructed an international network of foreign investments using the countries responsible for the fishing activities (vessel flags) and the shareholder’s country of origin. This network representation allowed us to explore the likelihood of investments (probability of a link) since investing in a particular country is not independent of network structure (e.g. investments already present in the network, or individual features of the countries involved). We used governance data from the World Bank (control of corruption, government effectiveness, political stability, regulatory quality, rule of law, and voice accountability) to investigate if foreign investments in fishing activities exposed to marine tipping points tend to occur in countries with lower effectiveness and weaker rule of law, as previously reported in the context of emerging zoonotic diseases27. To test this hypothesis, we used an exponential random graph model44 where nodes are countries and links are investment flows between shareholders and companies running fishing operations.
@@ Results
Figure 3: Actors exposed to marine tipping Top companies (A) and shareholders (B) fishing in areas with high risk of regime shifts. Colours correspond to country and countours to shareholder types (B). The rank for global ultimate owners is reported in Fig. S2
@@
Out of the 114,191 vessels reported in the Global Fish Watch database, we found that 0.15 % of all vessels tracked (N = 16,878) fish in areas prone to marine tipping points (Fig 1). The bulk of these vessels do not report a flag or country of registry; however, for those who do, the largest fleets are associated to China, the United States, Japan and Norway (Fig 2A, see Fig S1 for vessels types ). China’s fleet is at least three times larger than the next larger fleet, making it a dominant actor. Out of the 16,878 vessels, we identified 2,981 companies, for 855 of them we identified 1,825 unique shareholders and 2,783 ownership relationships between shareholders and companies. Most of the shareholders are based in Norway, China, Russia, Spain, the United Kingdom, and Denmark (Fig 2B); the most common shareholder types are corporate entities and individual families.
@@
The network of ownership and shareholder investments is relatively sparse, lacking the presence of major financial giants common to multiple companies. Unlike previous studies that identified “keystone actors” – central corporations and financial entities with disproportionate influence in the market15 and potential exposure to tipping elements of the Earth system26– our findings reveal a more diffuse network with no central actors. This is similar to the pattern found by Jouffray et al. when looking at shareholders of publicly listed seafood companies29. One possible explanation is the ownership structure where most fishing companies in our dataset are owned by only a few shareholders who in turn own on average half, if not all, of the company’s shares. (Fig. 2D, E). This is reflected by a large proportion of our shareholders being individuals, families, or small-scale businesses (Fig 2C). Another possibility is that keystone actors would emerge in the network only when considering not just vessel owners, but also retailers and traders involved in the supply chain.
@@
We found that government effectiveness was the best predictor of foreign investments (Fig 4). Countries with effective governments tended to invest in governments that have less effectiveness (hence the positive difference in the full model). We also found a smaller, yet significant and negative effect on political stability. Countries with low political stability in our data set invested in companies located in countries with slightly higher political stability (hence the negative difference on the full model).
Despite data incompleteness, our results show that if marine tipping points were to synchronize globally, they would impact an industry generating over 6.6 billion dollars in revenue and affect more than 100,000 jobs (Fig 5). The vessel flags producing the highest revenues in our data were Ecuador, Panama, Russia, Spain and the UK, while those employing the largest workforces were led by Russia, Japan, Ecuador, Panama and Argentina (Fig 5). Vessel flags are not necessarily where profits go, but data gaps prevented a more accurate approximation. We were able to retrieve revenue and employment data for only 2,216 companies, with roughly 70% of the data missing values. As a result, our figures likely underestimate the true potential economic and social impacts of possible declines in primary productivity from marine heatwaves and marine regime shifts. Much higher levels of transparency in company reporting are needed to improve these initial estimates.
Figure 5: Potential impacts of marine tipping points. Countries ranked by the revenues reported in our dataset (A). Only countries with revenues higher than 1M US\$ shown. Countries ranked by workforce employeed (B), only countries with more than 500 employees shown. The data however has high percentage of missing values ( 70%), so it should be interpreted as an under estimation.

### v2

@@ Introduction
Forecasting the ecological and economic consequences of marine regime shifts is a difficult endeavor due to uncertainty. Increasing sea surface temperature can destabilize upwellings – currents that bring nutrient-rich water to the ocean’s surface – and affect some of the world’s most productive fisheries11. While insufficient nutrients can reduce primary productivity and cascade up through marine food webs, an excess can cause coastal eutrophication and hypoxia6, with detrimental consequences to fish stocks. Climate change is expected to reduce fish biomass by 3 to 22% in tropical and temperate areas by the end of the century12. Overfishing and nutrient pollution from agriculture and urban waste can exacerbate this impact. A combination of oceanographic and climatic factors heightens the risk of regime shifts, putting both ecosystems and human livelihoods at risk2,13. Yet, observing and measuring tipping points is a challenging task typically masked by low signal to noise ratios14. However, proactive management strategies that account for these risks could help mitigate exposure to tipping points2.
@@
Companies and financial institutions are key actors that contribute to the extraction of marine resources, while also being affected economically by changes in marine ecosystems. About 35.4% of fish stocks globally are currently depleted or overexploited15, with only 13 companies controlling around 16% of wild seafood capture16,17. High market concentration, a form of inequality, means that a few actors can have disproportional effects on the system for good or bad18. Fishing companies contribute to ecological degradation through harmful practices like overfishing19, destructive fishing methods20, and non-compliance with regulations21,22. By removing biomass from the ocean, fishing has historically prevented the sequestration of 21.8M tons of carbon15,23. In addition, the industry’s return on investment is becoming increasingly risky. Modern fishing fleets travel twice as far but catch only one-third of what they used to per kilometre travelled24, it is estimated that 54% of high seas fishing would be unprofitable without public subsidies15,25. Companies doing extractive activities other than fisheries are also projected to have impacts from climate change from extreme weather events and sea level rise26,27
@@
Figure 1: Areas at risk of tipping points. We computed extreme and severe heatwave events based on historical records of sea surface temperature (A, B). Areas in colour are places where the probability is non-zero, expressed as the mean number of days spent on a heatwave event. Ocean colour data was used to compute early warning signals of critical transitions (C). Note that there are no warnings computed for higher latitudes due to missing data. For the analysis of areas prone to marine regime shifts, we kept areas where the probability of extreme and severe heatwaves is non-zero, or where at least three early warning signals were found (see Methods). Supplementary figure S1 provides additional empirical evidence of the occurrence of regime shifts by mapping 1448 case studies coded through literature review.
@@
It should be in the best interest of companies and their shareholders to act as stewards of the seas and maintain a productive and healthy ocean28,29. Financial actors – such as banks, asset managers, and pension funds – could play an important role in conservation by deciding not to finance economic activities that alter the ecological integrity of the biosphere30–33. For instance, a small set of financial actors have an increasing influence over climate stability through their equity holdings in companies operating in the Amazon and boreal forests30. Likewise, large asset managers and other financial institutions invest in companies operating in areas vulnerable to the emergence of infectious diseases31. This leaves them exposed to the impacts of their investments34.
Figure 2: Abrupt changes in primary productivity. We used the structural change and Pettitt test to investigate abrupt shifts in Chlorophyll A data and computed the proportion of change before and after the shift (the mean after over the mean before). If the proportion is one there was no change in means, if the proportion is < 1 it means an abrupt decline (e.g. 0.5 is a 50% decline), or an increase if >1 (A). 16% of the oceans show significant abrupt changes with the structural change test (B) while 50% of the oceans show abrupt changes in primary productivity with the Pettitt test (C). Supplementary Figure S2 shows a similar test for linear trends in primary productivity.
@@
Climate-change-driven extreme weather events are expected to cost up to USD 14 trillion annually by 210035, directly threatening the solvency of insurers and reinsurers36. Investors and asset managers face climate exposure being at risk of stranded assets, changes in commodity prices, scarcity of resources, disruption of supply chains, and infrastructure damages37. In response, financial actors are increasingly expected to disclose their risks through initiatives such as the Task Forces on Climate- and Nature-Related Financial Disclosures (TCFD and TNFD, respectively) or the EU non-financial reporting directive.
Figure 3: Abrupt changes in fish biomass. We tested abrupt changes in fish biomass data from the RAM legacy database using the Pettitt and structural change tests. We found 252 stocks having abrupt changes detected by both tests, 14 only by the structural change F statistic, 83 only by the Pettitt test, and 20 with no detection (A). The mean decline for all significant changes was 50% of the biomass with respect to the mean before the abrupt change (B). The bottom panel lists the top territories most affected (C), the geographic areas most impacted (D), and the species with most abrupt changes. Supplementary Figure S3 complements the analysis by comparing the timing of the shifts between tests and a linear test.
@@
Analyses of financial exposure to biodiversity loss and ecosystem degradation are mostly focused on terrestrial ecosystems, with little research addressing ocean-related financial risks38. Yet, fishing companies, their shareholders, and the countries they operate in are exposed to potentially irreversible changes and consequent declines in ecosystem services, with consequences on the economy, investments and livelihoods. However, being exposed to regime shifts also means being in a position to take actions that reduce anthropogenic pressures that could drive the system across a potential tipping point. In this paper, we asked who are these corporate and financial actors exposed to marine tipping points?
@@ Methods
Areas at risk: We identified areas of the world vulnerable to marine regime shifts. We approximated these areas as places currently under the strong influence of marine heatwaves or already showing symptoms of resilience loss. Frölicher et al.8 defined heatwaves as events when sea surface temperature exceeds the 99th percentile with respect to daily satellite observations for a 30-year period. Hobday et al.39 further classified heatwave categories based on the 90th percentile, where moderate heatwaves are between 1x and 2x the 90th percentile, strong heatwaves between 2x and 3x, severe heatwaves between 3x and 4x, and extreme heatwaves above 4 times the 90th percentile. We computed the probability of severe and extreme heat waves based on daily sea surface temperature observations from the US National Oceanic and Atmospheric Administration (NOAA) using the optimum interpolation data from 1981 to 2022 at 0.25 resolution (Fig 1A-B). While several impacts of heatwaves have been documented in ecosystems and the economy10,39, a recent global assessment showed no systematic effect on fish biomass40. However, heatwaves by themselves are seldom the sole cause of fish populations decline, it is often the compound effect of temperature with other pressures such as low primary productivity or low oxygen zones13.
@@
Thus we complemented heat waves detection with proxies of stability loss in primary productivity, which are in turn driven by oxygen depletion, salinity, and the mixing of the upper layer of the ocean6,7,41. We used a geographically explicit mapping of marine ecosystems showing symptoms of resilience loss42. Resilience is the ability of a system to withstand disturbances without losing its functions and identity43. When systems lose resilience and are close to tipping points, they leave statistical signatures in time series known as early warning signals44. The map was produced by computing temporal autocorrelation, standard deviation, skewness, kurtosis and fractal dimension on Chlorophyll-A concentration, a proxy of marine primary productivity, using data from the European Space Agency (weekly observations at 0.25 resolution from 1998-2018)45. The data are limited by the coverage of the ocean colour data (the 50 N-50 S region), excluding the Arctic and Antarctic oceans42. For this paper we used areas with at least 3 early warnings, since two of them are suitable for bifurcation tipping (autocorrelation and standard deviation), two of them are suitable for flickering and stochastic transitions (skewness and kurtosis), and the fractal dimension is robust against heteroskedasticity. At least three early warnings give us the chance to pick up more robust signals of the risk of tipping taking into account different ways of reaching the transition.
Evidence of regime shifts: We documented past regime shifts by reviewing scientific literature using the framework of the regime shifts database46. In addition, we performed change point analysis, a classical method to detect regime shifts in time series47,48. We looked for break points on the Chlorophyll-A concentration data from the biological pump and carbon exchange process project49 which offers monthly data at 9km pixel resolution, and is derived from the European Space Agency’s Ocean Colour Climate Change initiative. We also looked at break points on historical fish stock assessment data from the RAM legacy stock assessment database version 4.4450,51, similar to previous efforts identifying regime shifts on fish stocks48. Out of the 479 stocks for which there are time series reported, we assessed only stocks with more than 25 years with data (N = 373).  If there is an abrupt decline in primary productivity, one can expect that the foodweb depending on it can be prone to similar changes. However, linear changes in primary productivity can also generate abrupt decline on fish stocks. For that reason we also calculated the linear trend on Chlorophyll-A time series to show areas of the world with significant changes by using the Kendall tau statistic. For break point detection we used the Pettitt test of change in means52, and the change point F test statistic for change on residuals53.
@@
Fishing pressure: We identified fishing vessels operating in pixels showing a high probability of heatwaves or symptoms of resilience loss. We used the dataset published by Kroodsma et al.54 who identified >70,000 industrial fishing vessels in the size range from 6 to 146m in length, thanks to an automatic identification system (AIS). Vessels need to broadcast their position on the AIS to avoid potential collisions, so the data offers individual vessel trajectories from 2012-2020 (fishing effort dataset from Global Fish Watch version 2.0). We extracted vessels whose fishing activities were recorded in areas with high probabilities of heatwaves (severe or extreme events) or high probability of regime shifts. We only include pixels with at least three different early warning signals of resilience loss, that way we include multiple routes of tipping (e.g. noise-induced, flickering, not only the traditional slowing down), and keep pixels for which multiple lines of evidence suggest loss of resilience.
@@
Owners and shareholders: We identified the individual and company beneficial owners of the vessels using the maritime mobile service identity to match the Global Fishing Watch data with vessel owners in the Lloyds database (www.seasearcher.com). We matched 13,537 records, but not all of them had available ownership information. We also used the dataset curated by Carmine et al.55 where vessels fishing on the high seas were manually annotated. We added companies whose vessels matched our selection of vessels fishing in areas with marine tipping point risk. To identify shareholders and global ultimate owners of the companies responsible for fishing activities, we used Orbis, a database of companies around the world with coverage of > 489 million companies. Where possible, we extracted data about their shareholder structure (e.g. direct and total ownership, shareholder type), annual revenues, number of employees, and shareholder’s countries of operation. Unfortunately, Orbis and Lloyds are not open access data sources and their terms of service forbids sharing the raw accessed data with third parties. Both of them were accessed through research subscriptions by the home institution.
@@
Network models: We constructed an international network of foreign investments using the countries responsible for the fishing activities (vessel flags) and the shareholder’s country of origin. This network representation allowed us to explore the likelihood of investments (probability of a link) since investing in a particular country is not independent of network structure (e.g. investments already present in the network, or individual features of the countries involved). We used governance data from the World Bank (control of corruption, government effectiveness, political stability, regulatory quality, rule of law, and voice accountability) to investigate if foreign investments in fishing activities exposed to marine tipping points tend to occur in countries with lower effectiveness and weaker rule of law, as previously reported in the context of emerging zoonotic diseases31. To test this hypothesis, we used an exponential random graph model56 where nodes are countries and links are investment flows between shareholders and companies running fishing operations.
@@ Results
Figure 5: Actors exposed to marine tipping Top companies (A) and shareholders (B) exposed to areas with high risk of regime shifts. Colours correspond to country and countours to shareholder types (B). The rank for global ultimate owners is reported in Fig. S5
@@
We found that 30% of global oceans have presented extreme heat waves, and 66% severe events (Fig 1). Most of the extreme events are located in high latitudes in the Arctic and Antarctic oceans. Severe events are also common in the tropical Pacific ocean (related to ENSO), coastal California, and the Indian ocean. Upwelling systems in the coast of California, Benguela or Canary islands show at least three early warning signals of the proximity of instabilities. 24% of the area analysed with early warning signals show symptoms of resilience loss42. The tropical north Atlantic is also a hotspot of resilience loss, a place where previous work has documented increase in summer time stratification of the water column driven by sea surface temperature and salinity41. With stronger stratification there is less mixing of water column and nutrient cycling, leading to a decline in primary productivity. We detect such signal in the linear decline of primary productivity (Fig S2).
Our change point analysis shows that there has been abrupt changes in primary productivity over time (Fig 2). The extend of the area varies depending on the statistical test used. The Pettitt test looks for changes in means, and detects significant changes in 50% of the oceans with break points in early 2000s in the south Pacific, and more recent changes in the rest of the world. The structural change test is more restrictive, seeking changes on the residuals of a lagged autoregressive model. We find again abrupt changes in the central Pacific, the tropical Atlantic, the coasts of California, and India; adding up to 16% of the oceans. For the test that looks if there is a decline on Chlorophyll-A regardless if it is abrupt or smooth (the slope of a linear regression), we found that 77.3% of the ocean area (50N-50S region) has a significant kendall tau, and for 41.9% of the area primary productivity is declining (Fig S2).
Similarly, for the stock assessment data we found that 353 fish stocks show significant abrupt changes of which 271 are declining biomass while 82 are increasing (Fig 3). We find agreement between the two methods used for change point detection (256 stocks), but some differences in the time predicted. The mean decline between the mean before and the mean after the break point is of 50% of the biomass. Qualitatively, our results align with previous studies using different methods, smaller sample sizes but stricter thresholds for collapse48,57–59, where the proportion of stocks collapsed are between 25% and 50%. Here we find that 119 stocks collapsed below 50% their historical average (31%), while 29 stocks (7%) collapsed below 15% their historical record (Fig 3B). Regardless if the changes are abrupt, the Sen slope test supports the direction of change identified through the change point analysis (Fig S2). The countries with more fish stocks showing abrupt changes are USA, Canada and Australia, although many occur in international waters; the fishing grounds where most abrupt changes reported are the Pacific coast, the Gulf of Alaska, central and southern Chile and the Gulf of Mexico. The species most impacted are herring, Atlantic cod, walley pollock, red rock lobster and pacific herring (Fig 3).
@@
Out of the 114,191 vessels reported in the Global Fish Watch database, we found that 15 % of all vessels tracked (N = 16,878) fish in areas prone to marine tipping points (Fig 1). The bulk of these vessels do not report a flag or country of registry; however, for those who do, the largest fleets are associated to China, the United States, Japan and Norway (Fig 4A, see Fig S4 for vessels types ). China’s fleet is at least three times larger than the next larger fleet, making it a dominant actor. Out of the 16,878 vessels, we identified 2,981 companies, for 855 of them we identified 1,825 unique shareholders and 2,783 ownership relationships between shareholders and companies. Most of the shareholders are based in Norway, China, Russia, Spain, the United Kingdom, and Denmark (Fig 4B); the most common shareholder types are corporate entities and individual families.
@@
The network of ownership and shareholder investments is relatively sparse, lacking the presence of major financial giants common to multiple companies. Unlike previous studies that identified “keystone actors” – central corporations and financial entities with disproportionate influence in the market16 and potential exposure to tipping elements of the Earth system30– our findings reveal a more diffuse network with no central actors. This is similar to the pattern found by Jouffray et al. when looking at shareholders of publicly listed seafood companies33. One possible explanation is the ownership structure where most fishing companies in our dataset are owned by only a few shareholders who in turn own on average half, if not all, of the company’s shares. (Fig. 4D, E). This is reflected by a large proportion of our shareholders being individuals, families, or small-scale businesses (Fig 4C). Another possibility is that keystone actors would emerge in the network only when considering not just vessel owners, but also retailers and traders involved in the supply chain.
@@
We found that government effectiveness was the best predictor of foreign investments (Fig 6). Countries with effective governments tended to invest in governments that have less effectiveness (hence the positive difference in the full model). We also found a smaller, yet significant and negative effect on political stability. Countries with low political stability in our data set invested in companies located in countries with slightly higher political stability (hence the negative difference on the full model). The implications of this result is that the relevant decisions to manage tipping point risk are not only in the hands of fishing vessels, the companies that own them or their shareholders. The countries where these activities take place (fishing and investing) can design regulations that restrain or disincentivise investments in other regions of the world where fishing activities are happening in high risk areas.
Despite data incompleteness, our results show that if marine tipping points were to synchronize globally, they could impact an industry generating over 6.6 billion dollars in revenue and affect more than 100,000 jobs (Fig 7). The vessel flags producing the highest revenues in our data were Ecuador, Panama, Russia, Spain and the UK, while those employing the largest workforces were led by Russia, Japan, Ecuador, Panama and Argentina (Fig 7). Vessel flags are not necessarily where profits go, but data gaps prevented a more accurate approximation. We were able to retrieve revenue and employment data for only 2,216 companies, with roughly 70% of the data missing values. As a result, our figures likely underestimate the true potential economic and social impacts of possible declines in primary productivity from marine heatwaves and marine regime shifts. Much higher levels of transparency in company reporting are needed to improve these initial estimates.
Figure 7: Potential impacts of marine tipping points. Countries ranked by the revenues reported in our dataset (A). Only countries with revenues higher than 1M US\$ shown. Countries ranked by workforce employeed (B), only countries with more than 500 employees shown. The data however has high percentage of missing values (7̃0%), so it should be interpreted as an under estimation.